# Hawking Radiation Signatures from Primordial Black Holes Transiting the Inner Solar System: Prospects for Detection

### v1

@@ II.3.1 Dark Matter Density Profile
$\begin{split}&\rho_{0}=0.0196\,M_{\odot}\text{pc}^{-3},\\ &r_{0}=15.5\,\text{kpc},\\ &r_{vir}=287,\\ &q=1.22.\\ \end{split}$ (33)
@@ II.3.2 Dark Matter Velocity Profile
Figure 7: Maxwellian velocity profile for dark matter located $8$ kpc from the center of the Milky Way. $v_{\odot}$ is azimuthal velocity of the Sun in the galactic plane and $v_{\text{rms}}$ is the root mean square velocity for the Maxwellian distribution. The distribution drops to zero at the galactic escape velocity $v_{\rm esc}\approx 544{\rm km/s}$.
@@ II.4.1 Geomagnetic Field
The geomagnetic field can be approximated by the $l=1$ dipole term of Eq. (37) with magnetic moment $7.7\times 10^{22}$m2 Alken et al. (2021). The magnetic moment is not static spatially or temporally and has fluctuated and reversed over the course of Earth’s history. Recent measurements show the dipole moment has been decreasing since at least the year 1600 Korte and Muscheler (2012). Archeomagnetic and paleomagnetic data from core samples, thermoremnant magnetization, and archaeological samples show complex variation of the dipole moment over large time-scales Korte and Muscheler (2012); Panovska et al. (2019), and evidence that the geomagnetic field reverses over uncorrelated intervals of tens of thousands to millions of years in a statistically random process Gaffin (1989); Phillips and Cox (1976). The dipole axis is not parallel to the Earth’s rotation axis, and the magnetic north dipole is currently located near $86.5^{o}$ latitude Alken et al. (2021). The locations of the magnetic poles drift with time, and are projected by the most recent International Geomagnetic Reference Field report to travel at an average rate of $40$km/yr from 2020-2025 Alken et al. (2021). Since 1900, the magnetic North pole has traveled northward by approximately $15^{o}$ latitude Alken et al. (2021).
@@
This cutoff is the minimum rigidity (momentum per unit charge) necessary for a particle to reach a point in the magnetosphere from infinity, and it is generally a function of particle path, location, and geomagnetic field Kress et al. (2015). $R_{c}$, with units of MV, is a function of radius $r$ in cm, magnetic latitude $\theta$, azimuthal angle $\xi$, zenith angle $\epsilon$, and magnetic moment $\mathcal{M}$ in $\text{G}\cdot\text{cm}^{3}$, all defined in Ref. Smart and Shea (2005). The geomagnetic cutoff rigidity ranges from about $15$ GV at the equator to 00 GV at the magnetic poles, where the tangential component of the field vanishes. Sub-GeV charged cosmic rays can therefore only be detected by an experiment in LEO when its orbit takes it near the magnetic poles—where charged particles can penetrate to lower altitudes. This reduces both the measured positron signal and background and significantly reduces the proton background at low energies for an experiment in LEO. We include the effect of the geomagnetic cutoff in our simulations of time-dependent positron signals from PBH transits measured by AMS in section III.2.
@@ III.1 PBH Transit Rates
[…]
@@
Figure 9: Number of expected positron-producing PBHs that pass within $1\,{\rm AU}$ of Earth per year. Here we plot the yearly transit rate $\bar{\Phi}_{e^{\pm}}$ according to Eq. (41) with $b_{\rm max}=1\,{\rm AU}$, $t=t_{0}$, and $f_{\rm PBH}=1$, as a function of the mass $\bar{M}_{i}$ at the peak of the formation-time PBH mass function, given in Eqs. (11) and (13) for the GCC and LN functions, respectively. Rates for the LN mass function range from $0.005-1065$ yr-1 for $0.1\leq\sigma\leq 2.0$. The GCC PBH transit rate (top right) is plotted as a function of $\bar{M}_{i}$ and shown for three sets of parameters $\{\alpha,\beta\}$. Contour plots of the GCC PBH transit rate with $\alpha$ constant (bottom, left) and $\beta$ constant (bottom, right) indicate that the transit rate is only weakly sensitive to $\alpha$ and $\beta$ for $\bar{M}_{i}>M_{*}$. For both models, if $\bar{M}_{i}$ falls in the positron-producing mass range, $M_{*}\leq\bar{M}_{i}\leq 5\times 10^{17}$g, the expected transit rates exceed $1\,{\rm yr}^{-1}$, which implies that a decade of precision time-series data could probe large regions of parameter space. Vertical dashed lines at $\bar{M}_{i}=2\times 10^{17}\,{\rm g}$ indicate the mass where current evaporation constraints on extended mass functions stop and $f_{\rm PBH}=1$ is viable De la Torre Luque et al. (2024); Gorton and Green (2024). Note that detectable positron signals could arise even for $\bar{M}_{i}>5\times 10^{17}\,{\rm g}$ because we have incorporated realistic, extended mass functions and evolved them from the time of PBH formation to today; hence a subpopulation of PBHs with $M(t_{0})\leq 5\times 10^{17}\,{\rm g}$, capable of emitting positrons today, would emerge from the small-mass tail of the underlying PBH mass function.
@@
Given how modest the mass-loss due to Hawking evaporation is over the current age of the universe for $M_{i}=5\times 10^{17}\,{\rm g}\gg M_{*}=5.34\times 10^{14}\,{\rm g}$, we incur an error of only $(M(t_{0},M_{i})-M_{i})/M_{i}=-3.6\times 10^{-10}$ for $M_{i}=5\times 10^{17}\,{\rm g}$, per Eq. (19). Hence we may use the positron-emitting mass $M(t_{0})\simeq M_{i}=5\times 10^{17}\,{\rm g}$ when evaluating $\bar{\Phi}_{e^{\pm}}(b_{\rm max})$.
@@ III.2 Time-Dependent PBH Signature
Figure 10: Geometry for simulating the time-dependent $e^{+}$ signal from a PBH transit measured by a detector in orbit about the Earth. The blue field of view cone rotates with the detector area $\vec{A}(t)$ about a fixed point at frequency $\omega=2\pi/\tau$. We approximate the orbit as rotation about a fixed point because $𝐿b\gg R_{\mathTerra}+h$. The PBH (yellow circle) transits past on a linear trajectory with velocity $\bar{v}=246\,{\rm km/s}$ while emitting positrons at a rate $\dot{N}(M)$ given by Eq. (43).
@@ III.2.1 Simulated PBH Signals
We first compute the time-dependent positron signal generated by a transiting PBH and measured by an ideal detector in orbit around the Earth. The orbit, which approximates that of the ISS, is parameterized by a height $h=450$ km, a period $\tau=90$ min, and an inclination of ±$\pm 51.6^{o}$ latitude. Our ideal detector, which measures positrons with $100\%$ efficiency, has a field of view (FOV) angle $\theta_{\text{FOV}}$, an oriented area $\vec{A}(t)$, an energy-dependent acceptance $a(Q)$, and a minimum detectable positron energy $Q_{\rm min}$. This $Q_{\text{min}}$ parameter is determined in practice by the detector geometry, efficiency, trigger, and orbit relative to the geomagnetic field. For AMS, we take $Q_{\rm min}=500\,{\rm MeV}$, as discussed in Section II.4.2.
@@
[…]
@@
[…]
The weighting by $\tau_{\rm obs}$ in Eq. (43) approximately accounts for the changing geomagnetic cutoff due to variations in latitude during the orbits. Estimating observation time for positrons in a given energy bin is quite complicated for a real detector like AMS, which not only orbits the Earth and traverses ±$\pm 51.6^{o}$ of latitude, but is subject to various ISS maneuvers which change its pointing dynamically with time. Furthermore, the geomagnetic field is dynamic and has many features not captured by Stoermer’s dipole approximation, such as the South-Atlantic Anomaly, where field intensity drops to about 1/3 of the typical intensity and particle fluxes increase dramatically Pavón-Carrasco and De Santis (2016). A more accurate determination of observation time as a function of positron energy would require ISS trajectory and altitude data, AMS orientation data (which is collected by an onboard GPS), and the International Geomagnetic Reference Field—a precise model of the time-varying geomagnetic field Alken et al. (2021). This is beyond the scope of this analysis.
@@ III.2.2 Positron Background
For each PBH transit, positron fluxes at different energies are sampled from distinct Gaussian distributions, converted to particle counts, and summed together to form an energy-integrated background for the energy range $[Q_{\rm min},\,Q_{\rm max}]$. For the $i$th energy bin of width $\Delta E_{i}$ and characteristic energy $\tilde{E}_{i}$, we calculate the mean positron background count rate as measured by our detector, $\langle\zeta_{\text{bkgd},i}\rangle$ with units ${\rm s}^{-1}$:
@@
Therefore the SNR, and thereby the PBH detection efficiency, is determined by the variances of all the bins enclosed in the interval $[Q_{\rm min},Q_{\rm max}]$.
@@
To generate a daily flux background for each energy bin, we therefore sample from a Gaussian with mean $\langle\zeta_{\text{bkgd},i}\rangle$, calculated according to Eq. (49), and standard deviation $\sigma_{i}$. We then convert the fluxes to counts and sum over all energy bins in the range $[Q_{\rm min},Q_{\rm max}]$. An example of a sampled background superposed with the simulated PBH signal is shown in Fig. 14 (green points).
Figure 13: AMS positron daily fluxes collected over approximately 9 years of observation. The data displays a period of 11 years, corresponding to the 11 year solar cycle. Solar modulation effects are most prominent for low energy positrons. We plot the four reported lowest energy bins.
@@ III.2.4 Maximum Impact Parameters
Given that we can use numerical studies to determine detection efficiency, as plotted for example parameters $M=10^{14}\,{\rm g},\,b=0.15\,{\rm AU}$ in Fig. 15, we may next determine a maximum impact parameter function, $b_{\text{max}}(M)$, for each detector configuration. We define $b_{\rm max}$ as the largest impact parameter such that a PBH with mass $M(t_{0})$ can be detected by a given detector geometry at $5\sigma$ significance with 99% efficiency. The general steps followed to calculate $b_{\text{max}}$ for a specified value of $M(t_{0})$ are: (1) simulate $\mathcal{O}(10^{4})$ days of background, (2) for a given value of $b$, generate the optimal filter ${\cal H}(t|M,b,v)$, (3) run the noise through the filter ${\cal H}(t)$ to generate a background peak amplitude distribution, (4) simulate 1000 PBH transits with the given $b$ value and determine filtered peak amplitude significance for each one, (5) compute detection efficiency (the fraction of the 1000 transits which are detected at or above the $5\sigma$ level), (6) repeat steps 2-4 for impact parameters in the range $[10^{-3}\,{\rm AU},10^{2}\,{\rm AU}]$, (7) identify $b_{\text{max}}(M)$ as the value of $b$ at which detection efficiency drops below 99%. Figure 15 plots detection efficiency as a function of impact parameter $b$ for PBH mass $M(t_{0})=10^{14}\,{\rm g}$ using two different significance thresholds and assuming the AMS detector configuration.

### v2

@@ II.3.1 Dark Matter Density Profile
[…]
@@ II.3.2 Dark Matter Velocity Profile
Figure 7: Maxwellian velocity profile for dark matter located 88 kpc from the center of the Milky Way. $v_{\odot}$ is azimuthal velocity of the Sun in the galactic plane and $v_{\text{rms}}$ is the root mean square velocity for the Maxwellian distribution. The distribution drops to zero at the galactic escape velocity $v_{\rm esc}\approx 544{\rm km/s}$.
@@ II.4.1 Geomagnetic Field
The geomagnetic field can be approximated by the $l=1$ dipole term of Eq. (37) with magnetic moment $7.7\times 10^{22}$Am2 Alken et al. (2021). The magnetic moment is not static spatially or temporally and has fluctuated and reversed over the course of Earth’s history. Recent measurements show the dipole moment has been decreasing since at least the year 1600 Korte and Muscheler (2012). Archeomagnetic and paleomagnetic data from core samples, thermoremnant magnetization, and archaeological samples show complex variation of the dipole moment over large time-scales Korte and Muscheler (2012); Panovska et al. (2019), and evidence that the geomagnetic field reverses over uncorrelated intervals of tens of thousands to millions of years in a statistically random process Gaffin (1989); Phillips and Cox (1976). The dipole axis is not parallel to the Earth’s rotation axis, and the magnetic north dipole is currently located near $86.5^{o}$ latitude Alken et al. (2021). The locations of the magnetic poles drift with time, and are projected by the most recent International Geomagnetic Reference Field report to travel at an average rate of 4040km/yr from 2020-2025 Alken et al. (2021). Since 1900, the magnetic North pole has traveled northward by approximately $15^{o}$ latitude Alken et al. (2021).
@@
This cutoff is the minimum rigidity (momentum per unit charge) necessary for a particle to reach a point in the magnetosphere from infinity, and it is generally a function of particle path, location, and geomagnetic field Kress et al. (2015). $R_{c}$, with units of MV, is a function of radius $r$ in cm, magnetic latitude $\theta$, azimuthal angle $\xi$, zenith angle $\epsilon$, and magnetic moment $\mathcal{M}$ in $\text{G}\cdot\text{cm}^{3}$, all defined in Ref. Smart and Shea (2005). The geomagnetic cutoff rigidity ranges from about $15$ GV at the equator to 0 GV at the magnetic poles, where the tangential component of the field vanishes. Sub-GeV charged cosmic rays can therefore only be detected by an experiment in LEO when its orbit takes it near the magnetic poles—where charged particles can penetrate to lower altitudes. This reduces both the measured positron signal and background and significantly reduces the proton background at low energies for an experiment in LEO. We include the effect of the geomagnetic cutoff in our simulations of time-dependent positron signals from PBH transits measured by AMS in section III.2.
@@ III.1 PBH Transit Rates
$\begin{split}\bar{\Phi}_{\rm transit}(M_{\rm max},b_{\text{max}})&=\rho_{\rm DM}^{\odot}\int_{0}^{M_{\rm max}}dM\,\frac{1}{M}\psi(M,t_{0})\\ &\ \ \ \ \times\int_{0}^{v_{\rm esc}}dv\,vf(v)\int_{0}^{b_{\rm max}(M)}db\,2\pi b,\end{split}$ (40)
@@
Figure 9: Number of expected positron-producing PBHs that pass within $1\,{\rm AU}$ of Earth per year. Here we plot the yearly transit rate $\bar{\Phi}_{e^{\pm}}$ according to Eq. (41) with $b_{\rm max}=1\,{\rm AU}$, $t=t_{0}$, and $f_{\rm PBH}=1$, as a function of the mass $\bar{M}_{i}$ at the peak of the formation-time PBH mass function, given in Eqs. (11) and (13) for the GCC and LN functions, respectively. Rates for the LN mass function range from 0.00510650.005-1065 yr-1 for $0.1\leq\sigma\leq 2.0$. The GCC PBH transit rate (top right) is plotted as a function of $\bar{M}_{i}$ and shown for three sets of parameters $\{\alpha,\beta\}$. Contour plots of the GCC PBH transit rate with $\alpha$ constant (bottom, left) and $\beta$ constant (bottom, right) indicate that the transit rate is only weakly sensitive to $\alpha$ and $\beta$ for $\bar{M}_{i}>M_{*}$. For both models, if $\bar{M}_{i}$ falls in the positron-producing mass range, $M_{*}\leq\bar{M}_{i}\leq 5\times 10^{17}$g, the expected transit rates exceed $1\,{\rm yr}^{-1}$, which implies that a decade of precision time-series data could probe large regions of parameter space. Vertical dashed lines at $\bar{M}_{i}=2\times 10^{17}\,{\rm g}$ indicate the mass where current evaporation constraints on extended mass functions stop and $f_{\rm PBH}=1$ is viable De la Torre Luque et al. (2024); Gorton and Green (2024). Note that detectable positron signals could arise even for $\bar{M}_{i}>5\times 10^{17}\,{\rm g}$ because we have incorporated realistic, extended mass functions and evolved them from the time of PBH formation to today; hence a subpopulation of PBHs with $M(t_{0})\leq 5\times 10^{17}\,{\rm g}$, capable of emitting positrons today, would emerge from the small-mass tail of the underlying PBH mass function.
@@
Given how modest the mass-loss due to Hawking evaporation is over the current age of the universe for $M_{i}=5\times 10^{17}\,{\rm g}\gg M_{*}=5.364\times 10^{14}\,{\rm g}$, we incur an error of only $(M(t_{0},M_{i})-M_{i})/M_{i}=-3.6\times 10^{-10}$ for $M_{i}=5\times 10^{17}\,{\rm g}$, per Eq. (19). Hence we may use the positron-emitting mass $M(t_{0})\simeq M_{i}=5\times 10^{17}\,{\rm g}$ when evaluating $\bar{\Phi}_{e^{\pm}}(b_{\rm max})$.
@@ III.2 Time-Dependent PBH Signature
[…]
@@ III.2.1 Simulated PBH Signals
We first compute the time-dependent positron signal generated by a transiting PBH and measured by an ideal detector in orbit around the Earth. The orbit, which approximates that of the ISS, is parameterized by a height $h=450$ km, a period $\tau=90$ min, and an inclination of $\pm 51.6^{o}$ latitude. Our ideal detector, which measures positrons with $100\%$ efficiency, has a field of view (FOV) angle $\theta_{\text{FOV}}$, an oriented area $\vec{A}(t)$, an energy-dependent acceptance $a(Q)$, and a minimum detectable positron energy $Q_{\rm min}$. This $Q_{\text{min}}$ parameter is determined in practice by the detector geometry, efficiency, trigger, and orbit relative to the geomagnetic field. For AMS, we take $Q_{\rm min}=500\,{\rm MeV}$, as discussed in Section II.4.2.
@@
$\int_{0}^{Q_{\rm max}}\frac{d^{2}N^{(2)}}{dtdQ}dQ\bigg/\int_{0}^{\infty}\frac{d^{2}N^{(2)}}{dtdQ}dQ\geq 0.99.$ (42)
@@
$\dot{N}(M)=\int_{Q_{\text{min}}}^{Q_{\rm max}}dQ\,\tau_{\rm obs}(Q)\frac{d^{2}N_{e^{+}}^{(2)}}{dtdQ}(M,Q).$ (43)
The weighting by $\tau_{\rm obs}$ in Eq. (43) approximately accounts for the changing geomagnetic cutoff due to variations in latitude during the orbits. Estimating observation time for positrons in a given energy bin is quite complicated for a real detector like AMS, which not only orbits the Earth and traverses $\pm 51.6^{o}$ of latitude, but is subject to various ISS maneuvers which change its pointing dynamically with time. Furthermore, the geomagnetic field is dynamic and has many features not captured by Stoermer’s dipole approximation, such as the South-Atlantic Anomaly, where field intensity drops to about 1/3 of the typical intensity and particle fluxes increase dramatically Pavón-Carrasco and De Santis (2016). A more accurate determination of observation time as a function of positron energy would require ISS trajectory and altitude data, AMS orientation data (which is collected by an onboard GPS), and the International Geomagnetic Reference Field—a precise model of the time-varying geomagnetic field Alken et al. (2021). This is beyond the scope of this analysis.
@@ III.2.2 Positron Background
For each PBH transit, positron fluxes at different energies are sampled from distinct Gaussian distributions, converted to particle counts, and summed together to form an energy-integrated background for the energy range $Q_{\rm min},\,Q_{\rm max}]$. For the $i$th energy bin of width $\Delta E_{i}$ and characteristic energy $\tilde{E}_{i}$, we calculate the mean positron background count rate as measured by our detector, $\langle\zeta_{\text{bkgd},i}\rangle$ with units ${\rm s}^{-1}$:
@@
Therefore the SNR, and thereby the PBH detection efficiency, is determined by the variances of all the bins enclosed in the interval $Q_{\rm min},Q_{\rm max}]$.
@@
To generate a daily flux background for each energy bin, we therefore sample from a Gaussian with mean $\langle\zeta_{\text{bkgd},i}\rangle$, calculated according to Eq. (49), and standard deviation $\sigma_{i}$. We then convert the fluxes to counts and sum over all energy bins in the range $Q_{\rm min},Q_{\rm max}]$. An example of a sampled background superposed with the simulated PBH signal is shown in Fig. 14 (green points).
Figure 13: AMS positron daily fluxes collected over approximately 9 years of observation. The data displays a period of 11 years, corresponding to the $1$ year solar cycle. Solar modulation effects are most prominent for low energy positrons. We plot the four reported lowest energy bins.
@@ III.2.4 Maximum Impact Parameters
Given that we can use numerical studies to determine detection efficiency, as plotted for example parameters $M=10^{14}\,{\rm g},\,b=0.15\,{\rm AU}$ in Fig. 15, we may next determine a maximum impact parameter function, $b_{\text{max}}(M)$, for each detector configuration. We define $b_{\rm max}$ as the largest impact parameter such that a PBH with mass $M(t_{0})$ can be detected by a given detector geometry at $5\sigma$ significance with 99% efficiency. The general steps followed to calculate $b_{\text{max}}$ for a specified value of $M(t_{0})$ are: (1) simulate $\mathcal{O}(10^{4})$ days of background, (2) for a given value of $b$, generate the optimal filter ${\cal H}(t|M,b,v)$, (3) run the noise through the filter ${\cal H}(t)$ to generate a background peak amplitude distribution, (4) simulate 1000 PBH transits with the given $b$ value and determine filtered peak amplitude significance for each one, (5) compute detection efficiency (the fraction of the 1000 transits which are detected at or above the $5\sigma$ level), (6) repeat steps 2-4 for impact parameters in the range $10^{-3}\,{\rm AU},10^{2}\,{\rm AU}]$, (7) identify $b_{\text{max}}(M)$ as the value of $b$ at which detection efficiency drops below 99%. Figure 15 plots detection efficiency as a function of impact parameter $b$ for PBH mass $M(t_{0})=10^{14}\,{\rm g}$ using two different significance thresholds and assuming the AMS detector configuration.